# Electronic and magnetic properties of graphene-fluorographene nanoribbons: controllable semiconductor-metal transition

### v1

@@ Abstract
We investigate the electronic and magnetic properties of graphene channels ($2$-$4$ nm wide) embedded within fluorographene, focusing on two distinct interfaces: the fully-fluorinated $\alpha$ interface and the half-fluorinated $\beta$ interface. Density Functional Theory (DFT) calculations reveal that $\alpha\alpha$ systems exhibit semiconducting behavior with antiferromagnetic ordering, closely resembling pristine zigzag graphene nanoribbons. In contrast, $\alpha\beta$ systems display ferromagnetism and a width-dependent semiconductor-to-metal transition. To enable the study of larger systems, we develop and validate effective Hubbard models for both $\alpha\alpha$ and $\alpha\beta$ channels. Building upon DFT results and a Wannier function analysis, these models accurately reproduce the electronic structure and magnetic ordering observed in DFT calculations. Crucially, our $\alpha\beta$ model successfully captures the semiconductor-to-metal transition. Application of this model to larger systems reveals the persistence of a ferromagnetic state with spin polarization localized at the $\alpha$ edge. Our results demonstrate the potential of fluorination for targeted property engineering and provide a basis for exploring graphene-fluorographene systems in device applications ranging from microelectronics to spintronics.
@@ I Introduction
The paper is organized as follows: Section II describes our DFT methodology. Sections III and IV present the crystalline and electronic structures of the zigzag graphene channels. In Section V, we use Wannier90 calculations to derive a tight-binding Hamiltonian, capturing the essential electronic states near the Fermi level. Section VI builds an AH model from these insights, reproducing key graphene nanoribbon properties. Section VII validates the AH model against DFT results for wider graphene channels, allowing us to confidently extend our studies to system sizes that are computationally prohibitive for DFT calculations. Finally, Section VIII provides a summary and future perspectives.
Figure 1: (Color online). Relaxed crystalline structures of the graphene nanoribbons with fluorographene interfaces. These systems are graphene channels carved on fluorographene. The graphene-fluorographene $\alpha$ interfaces are enclosed by dashed lines in panel (a). Panels (b) and (c) show the $10$-$\alpha\alpha$ unit cell along the $x$ longitudinal (b) and $y$ transversal directions (c). Green balls are F atoms.
@@ II Computational approach
We performed density functional theory (DFT) calculations using the Quantum Espresso (QE) package [23] to investigate the electronic and magnetic properties of graphene channels. The Perdew-Burke-Ernzerhoff (PBE) exchange-correlation functional [24] was employed, along with a plane-wave basis set for electronic wave functions and charge density. Energy cutoffs of $70$ Ry and $420$ Ry were applied, respectively. We used ultrasoft pseudopotentials [25] to describe ion-electron interactions. Brillouin zone (BZ) sampling was tailored to the system’s electronic structure: (i) semiconductors, uniform $24\times 1\times 1$ $k$-point mesh; (ii) metals, uniform $36\times 1\times 1$ $k$-point mesh; (iii) near semiconductor-metal transition, finer uniform $54\times 1\times 1$ $k$-point mesh. Gaussian smearing was applied, with a degauss value of $0.005$ Ry for systems away from the transition and $0.001$ Ry near the transition. Crystalline structures were relaxed until forces and stress on atoms were below $0.001$ eV/Å  and $0.5$ GPa, respectively.
@@ III Crystalline structure of the graphene channels
Figure 2: (Color online). Relaxed crystalline structures of $n$-$\alpha\beta$ graphene-fluorographene nanoribbons ($n=10$), similar to Fig. 1 but with $\alpha\beta$ interfaces. The graphene channel interfaces are labeled as $\alpha$ (fully fluorinated, left) and $\beta$ (half-fluorinated, right). The dashed lines in (a) highlight these contrasting interfaces.
Figures 1 and 2 illustrate these interfaces (highlighted by dashed lines and shading). In the $\alpha$ interface, all carbon atoms of the bordering graphene zigzag chain are fluorinated, while only alternating carbon atoms are fluorinated in the $\beta$ interface. Panels (b) and (c) show top and side views of the structures. The unit cell repetition, due to the periodic boundary conditions, generates an array of alternating graphene channels, which are separated by fluorographene regions.
@@ IV Electronic structure of graphene channels
Figure 3: (color online) Band diagram ($\mathrm{E_{k}}$) and local density of states (LDOS) for $6$-$\alpha\alpha$ (top panels) and $6$-$\alpha\beta$ (bottom panels) graphene nanoribbons. The arrowheads in the LDOS plots indicate the positions of the $\alpha$ and $\beta$ interfaces. The $6$-$\alpha\alpha$ channel exhibits a semiconducting band gap ($\Delta=0.47$ eV), while the $6$-$\alpha\beta$ channel is near a semiconductor-metal transition. Near the Fermi level, the LDOS is mainly concentrated in the graphene (G) regions and has negligible contribution in the middle of the fluorographene (FG) regions, highlighting the insulating nature of fluorographene and its role in suppressing transverse electronic excitations, and leading to $k_{y}$-independent band structures in this energy range.
@@ IV.1 $\alpha\alpha$ channels
Figure 4 presents the spin-polarized band structures of $n$-$\alpha\alpha$ graphene-fluorographene superlattices for $n=6,8$ and $10$. The presence of an indirect energy gap ($\Delta$) between the valence ($b_{v}$) and conduction ($b_{c}$) bands confirms their semiconducting nature. Crucially, the gap magnitude decreases with channel width ($n$), ranging from approximately $0.47$ eV for $n=6$ to $0.31$ eV for $n=10$, see Table 1. This width-dependent band gap highlights the potential of these systems for tunable electronic properties, relevant for various semiconductor applications [1, 2, 3, 4, 5, 6, 7, 8].
Figures 5(a) and 5(b) reveals the spin density polarization, $m(\bm{r})=\rho_{\uparrow}(\bm{r})-\rho_{\downarrow}(\bm{r})$, within the graphene channel of the $10$-$\alpha\alpha$ system. This polarization exhibits a globally antiferromagnetic (AF) order: antiferromagnetic along the transverse direction to the interface and ferromagnetic along the longitudinal direction. The system’s net magnetization ($M_{t}$) is zero, resulting in spin-degenerate bands. This ordering is clearly visualized in the top view [Fig. 5(b)].
@@
Table 1 summarizes key electronic, magnetic, and structural properties of $n$-$\alpha\alpha$ systems. The fluorographene regions enforce a fixed longitudinal lattice parameter $\mathrm{a_{x}}=2.52$, while the transversal parameter ($\mathrm{a_{y}}$) increases with channel width. Despite the presence of spin polarization, all systems exhibit zero global magnetization ($M_{t}=0$) due to the antiferromagnetic order. The absolute magnetization $M_{abs}$ (the integrated absolute value of $m(\bm{r})$) increases with $n$ while the energy gap ($\Delta$) decreases. While the edge states spread more as the channel becomes wider, in agreement with the reduction of the energy gap, the magnetization at the edges, $M_{\alpha}=\pm 0.30$ $\mu_{B}$, remains the same.
Figure 5: (Color online) Spin density polarization, $m(\bm{r})=\rho_{\uparrow}(\bm{r})-\rho_{\downarrow}(\bm{r})$, of the $10$-$\alpha\alpha$ graphene-fluorographene system. (a) Side view: red and blue lobules represent opposite spin polarizations. (b) Top view: notice the antiferromagnetic order along the transverse direction to the interfaces and ferromagnetic order along the longitudinal direction. The system is globally antiferromagnetic (AF) with higher spin polarization at the channel edges. (c, d) Integration of the local density of states (ILDOS) within the energy window ($\mathrm{V}$) defined in Fig. 4(c). These states, located below the Fermi energy, exhibit $\mathrm{p_{z}}$ character, with higher weight at the channel edges.
@@ IV.2 $\alpha\beta$ channels
Figure 6: (Color online) Spin-polarized band structure of $n$-$\alpha\beta$ graphene-fluorographene systems for $n=4,6,8$ and $10$. Red and blue dots represent opposite spin states ($\uparrow$ and $\downarrow$). The ferromagnetic ground state leads to a spin-split band structure around the Fermi level. The band structure also reveals a semiconductor-metal transition with increasing channel width. At $\mathrm{k}=0$, the energy gap between the conduction ($b_{c}$) and valence ($b_{v}$) bands decreases with increasing $n$.
Figure 7: (Color online) Three-dimensional plot of the spin density polarization, $m(\bm{r})$, of the $n$-$\alpha\beta$ systems for $n=6,8$ and $10$. (a) and (e) show top views and (b)-(d) side views. The spin polarization is higher at the channel edges and decreases towards the center, with red and blue lobules representing opposite spin polarizations. The weight at the $\alpha$-edge remains constant, while it weakens at the $\beta$-edge with increasing channel width ($n\geq 8$) and nearly vanishes for $n=10$ (d, e). Top views (a) and (e) reveal a ferromagnetic order along the longitudinal direction and an antiferromagnetic order along the transverse direction. Despite this, the net magnetization is non-zero, making the $n$-$\alpha\beta$ nanoribbons ferromagnetic.
Figure 8: (Color online). Integrated local density of states (ILDOS) within the energy window ($\mathrm{V}$) defined in Fig. 6 for $\alpha\beta$ graphene-fluorographene systems. Similar to the $\alpha\alpha$ systems (Figs. 5c, 5d), the $\alpha\beta$ states near the Fermi level exhibit $\mathrm{p_{z}}$ character, with their weight concentrated along the graphene channel.
@@
A comparison of the band structures of $n$-$\alpha\alpha$ (Fig. 4) and $n$-$\alpha\beta$ (Fig. 6) reveals that the most significant differences, without considering the spin-splitting, occur at $\mathrm{k}\sim 0$. This indicates that the states of the valence and conduction bands near $\mathrm{k}\sim 0$ are strongly influenced by the distinct chemical nature of the $\beta$ interface, the key structural difference between the two systems.
@@
Figure 8 shows the ILDOS within the energy window $\mathrm{V}$ marked in Fig. 6. For the $6$-$\alpha\beta$ nanoribbon (Fig. 8a), the ILDOS strongly resembles the spatial profile of the spin density polarization (Fig. 7b), indicating that the states within the window $\mathrm{V}$ are the primary contributors to the spin-polarized regions. For the wider $8$-$\alpha\beta$ and $10$-$\alpha\beta$ systems, however, a notable difference arises at the $\beta$ edge between the ILDOS (Figs. 8b, 8c) and the spin density polarization (Figs. 7c, 7d) due to the reduced spin polarization of that edge, in agreement with the evolution of the band structure near $\mathrm{k}\sim 0$ (Fig. 6) with increasing channel width. Table 2 summarizes the key properties of $n$-$\alpha\beta$ graphene-fluorographene nanoribbons.
@@ V Tight-binding model
Figure 9: (Color online). Electronic band structure and projected density of states (PDOS) for bulk graphene and fluorographene. (a) Graphene band structure with $\pi$ and $\sigma$ states labeled. (b) Graphene PDOS. (c) Fluorographene band structure. (d, e) Fluorographene PDOS projected onto F and C atoms. States labeled $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{XY}$ in (c) will be used to describe the electronic structure of $n$-$\alpha\alpha$ and $n$-$\alpha\beta$ nanoribbons in the following sections.
@@
To identify the orbital character of the hybridized states in the nanoribbons, we first analyze the electronic structure of bulk graphene and fluorographene, along with their projected densities of states (PDOS), as shown in Fig. 9. Bulk graphene exhibits characteristic $\pi$ states near $\mathrm{E_{F}}$ with $\mathrm{p_{z}}$ orbital character [28]. Additionally, $\sigma$ states lie further away from $\mathrm{E_{F}}$ (Figs. 9(a) and 9(b)). In fluorographene (Figs. 9(c)-9(e)), we identify three key sets of states: (i) the $\mathcal{A}$ states, which are above $\mathrm{E_{F}}$, have a dominant $\mathrm{p_{z}}$ character mixed with $\mathrm{s}$ states; (ii) the $\mathcal{XY}$ states, which are below $\mathrm{E_{F}}$ and have primarily $\mathrm{p_{xy}}=\mathrm{p_{x}}+\mathrm{p_{y}}$ character; and (iii) the $\mathcal{B}$ states, found below $-4$ eV, and having mixed $\mathrm{p_{z}}$ and $\mathrm{s}$ character.
Figure 10: (Color online) Band structure of the $6$-$\alpha\alpha$ (a-c,g,h) and $6$-$\alpha\beta$ (d-f,i,j) systems. The projection (weight) on some relevant orbitals is highlighted. The size of each ($\mathrm{k},\mathrm{E}$) point reflects the weight of the eigenstates. The bands are classified based on the dominant orbital character from the reference states in Fig. 9: $\pi$ states ($\mathrm{p}_{z}$ character) from graphene (a, d). $\mathcal{A}$ and $\mathcal{B}$ states ($\mathrm{p}_{z}$ character) mainly located in the fluorographene zones (b, c, e, f). $\mathcal{XY}$ states ($\mathrm{p_{x}+p_{y}}$ character) mainly located in the fluorographene zones (g-j). In the $6$-$\alpha\alpha$ system, the conduction band ($b_{c}$) near $\mathrm{k}\sim$0 exhibits a hybridization of $\pi$ and $\mathcal{A}$ states, while the valence band ($b_{v}$) is primarily composed of $\pi$ states. A similar behavior is observed for the $6$-$\alpha\beta$ system, where both $b_{c}$ and $b_{v}$ bands are dominated by $\pi$ states.
@@
    dominant $\mathrm{\pi}$ character: Figs. 10(a) and 10(d) highlight the dominance of $\mathrm{\pi}$ states ($\mathrm{p_{z}}$ orbitals from graphene carbon atoms) near $\mathrm{E_{F}}$, spanning the energy range [8[-8[ - 8 eV, $12$ eV]]]]. The valence ($b_{v}$) and conduction ($b_{c}$) bands are primarily composed of these $\mathrm{\pi}$ states but exhibit crossings with bands of other orbital character.
@@
    hybridization with $\mathcal{A}$ states: Figs. 10(a)-10(f) reveal hybridization between $\mathrm{\pi}$ and $\mathcal{A}$ states contributing to the conduction band, with dominant C-$\mathrm{p_{z}}$ character. This is evident from the states within [1[1[ 1 eV, $12$ eV]]]] exhibiting both $\mathrm{\pi}$ and $\mathcal{A}$ orbital character due to their presence in both the graphene and fluorographene regions. The spin splitting observed in the $n$-$\alpha\beta$ bands near $\mathrm{k}=0$ (particularly in the $2$-$6$ eV window) further supports this $\mathrm{\pi}$-$\mathcal{A}$ hybridization.
@@
    minimal role of $\mathcal{XY}$ and $\mathcal{B}$ states: the lower panels of Fig. 10 demonstrate negligible hybridization between $\mathrm{\pi}$ and $\mathcal{XY}$ states (dominant F-$p_{xy}$ character, located within [1[-1[ - 1 eV, $-8$ eV]]]]). The absence of simultaneous $\mathcal{XY}$ and $\mathrm{\pi}$ character within the $b_{v}$ band, along with the lack of spin polarization affecting $\mathcal{XY}$ states, establishes their minimal contribution to $b_{v}$. Similarly, the $\mathcal{B}$ states (dominant F-$\mathrm{p_{z}}$ character, within [4[-4[ - 4 eV, $-11$ eV]]]]) lie well below the valence and conduction bands, making their contribution negligible.
As a consequence, to construct a simplified electronic model reproducing the $b_{v}$ and $b_{c}$ bands, the hybridization between $\mathrm{\pi}$ and $\mathcal{A}$ states is the crucial ingredient.

### v2

@@ Abstract
We investigate the electronic and magnetic properties of graphene channels (22224444 nm wide) embedded within fluorographene, focusing on two distinct interfaces: the fully fluorinated $\alpha$ interface and the half-fluorinated $\beta$ interface. Density functional theory (DFT) calculations reveal that $\alpha\alpha$ systems exhibit semiconducting behavior with antiferromagnetic ordering, closely resembling pristine zigzag graphene nanoribbons. In contrast, $\alpha\beta$ systems display ferromagnetism and a width-dependent semiconductor-to-metal transition. To enable the study of larger systems, we develop and validate effective Hubbard models for both $\alpha\alpha$ and $\alpha\beta$ channels. Building upon DFT results and a Wannier function analysis, these models accurately reproduce the electronic structure and magnetic ordering observed in DFT calculations. Crucially, our $\alpha\beta$ model successfully captures the semiconductor-to-metal transition. Application of this model to larger systems reveals the persistence of a ferromagnetic state with spin polarization localized at the $\alpha$ edge. Our results demonstrate the potential of fluorination for targeted property engineering and provide a basis for exploring graphene-fluorographene systems in device applications ranging from microelectronics to spintronics.
@@ I Introduction
The paper is organized as follows: Section II describes our DFT methodology. Sections III and IV present the crystalline and electronic structures of the zigzag graphene channels. In Sec. V, we use Wannier90 calculations to derive a tight-binding Hamiltonian, capturing the essential electronic states near the Fermi level. Section VI builds an AH model from these insights, reproducing key graphene nanoribbon properties. Section VII validates the AH model against DFT results for wider graphene channels, allowing us to confidently extend our studies to system sizes that are computationally prohibitive for DFT calculations. Finally, Sec. VIII provides a summary and future perspectives.
Figure 1: Relaxed crystalline structures of the graphene nanoribbons with fluorographene interfaces. These systems are graphene channels carved on fluorographene. The graphene-fluorographene $\alpha$ interfaces are enclosed by dashed lines in (a). (b), (c) The $10$-$\alpha\alpha$ unit cell along the (b) $x$ longitudinal and (c) $y$ transversal directions. Green balls are F atoms.
@@ II Computational approach
We performed density functional theory (DFT) calculations using the Quantum Espresso (QE) package [23] to investigate the electronic and magnetic properties of graphene channels. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [24] was employed, along with a plane-wave basis set for electronic wave functions and charge density. Energy cutoffs of $70$ and $420$ Ry were applied, respectively. We used ultrasoft pseudopotentials [25] to describe ion-electron interactions. Brillouin zone (BZ) sampling was tailored to the system’s electronic structure: (i) semiconductors, uniform $24\times 1\times 1$ $k$-point mesh; (ii) metals, uniform $36\times 1\times 1$ $k$-point mesh; (iii) near semiconductor-metal transition, finer uniform $54\times 1\times 1$ $k$-point mesh. Gaussian smearing was applied, with a degauss value of $0.005$ Ry for systems away from the transition and $0.001$ Ry near the transition. Crystalline structures were relaxed until forces and stress on atoms were below $0.001$ eV/Å  and $0.5$ GPa, respectively.
@@ III Crystalline structure of the graphene channels
Figure 2: Relaxed crystalline structures of $n$-$\alpha\beta$ graphene-fluorographene nanoribbons ($n=10$), similar to Fig. 1 but with $\alpha\beta$ interfaces. The graphene channel interfaces are labeled as $\alpha$ (fully fluorinated, left) and $\beta$ (half-fluorinated, right). The dashed lines in (a) highlight these contrasting interfaces.
Figures 1 and 2 illustrate these interfaces (highlighted by shaded regions). The $\alpha$ interface exhibits no fluorination, all carbon atoms of the bordering graphene zigzag chain are not bonded to fluorine atoms. In contrast, the $\beta$ interface displays partial fluorination, with alternating carbon atoms along the bordering zigzag chain bonded to fluorine atoms. Figures 1(b) and 1(c), as well as Figs. 2(b) and 2(c), show top and side views of the structures. The unit cell repetition, due to the periodic boundary conditions, generates an array of alternating graphene channels, which are separated by fluorographene regions.
@@ IV Electronic structure of graphene channels
Figure 3: Band diagram ($\mathrm{E_{k}}$) and local density of states (LDOS) for (a)-(d) $6$-$\alpha\alpha$ and (e)-(h) $6$-$\alpha\beta$ graphene nanoribbons. The arrowheads in the LDOS plots indicate the positions of the $\alpha$ and $\beta$ interfaces. The $6$-$\alpha\alpha$ channel exhibits a semiconducting band gap ($\Delta=0.47$ eV), while the $6$-$\alpha\beta$ channel is near a semiconductor-metal transition. Near the Fermi level, the LDOS is mainly concentrated in the graphene (G) regions and has negligible contribution in the middle of the fluorographene (FG) regions, highlighting the insulating nature of fluorographene and its role in suppressing transverse electronic excitations, and leading to $k_{y}$-independent band structures in this energy range.
@@ IV.1 $\alpha\alpha$ channels
Figure 4 presents the spin-polarized band structures of $n$-$\alpha\alpha$ graphene-fluorographene superlattices for $n=6,8$, and $10$. The presence of an indirect energy gap ($\Delta$) between the valence ($b_{v}$) and conduction ($b_{c}$) bands confirms their semiconducting nature. Crucially, the gap magnitude decreases with channel width ($n$), ranging from approximately $0.47$ eV for $n=6$ to $0.31$ eV for $n=10$; see Table 1. This width-dependent band gap highlights the potential of these systems for tunable electronic properties, relevant for various semiconductor applications [1, 2, 3, 4, 5, 6, 7, 8].
Figures 5(a) and 5(b) reveal the spin density polarization, $m(\bm{r})=\rho_{\uparrow}(\bm{r})-\rho_{\downarrow}(\bm{r})$, within the graphene channel of the $10$-$\alpha\alpha$ system. This polarization exhibits a globally antiferromagnetic (AF) order: antiferromagnetic along the transverse direction to the interface and ferromagnetic along the longitudinal direction. The system’s net magnetization ($M_{t}$) is zero, resulting in spin-degenerate bands. This ordering is clearly visualized in the top view [Fig. 5(b)].
@@
Table 1 summarizes key electronic, magnetic, and structural properties of $n$-$\alpha\alpha$ systems. The fluorographene regions enforce a fixed longitudinal lattice parameter $\mathrm{a_{x}}=2.52$, while the transversal parameter ($\mathrm{a_{y}}$) increases with channel width. Despite the presence of spin polarization, all systems exhibit zero global magnetization ($M_{t}=0$) due to the antiferromagnetic order. The absolute magnetization $M_{abs}$ [the integrated absolute value of $m(\bm{r})$] increases with $n$, while the energy gap ($\Delta$) decreases. While the edge states spread more as the channel becomes wider, in agreement with the reduction of the energy gap, the magnetization at the edges, $M_{\alpha}=\pm 0.30$ $\mu_{B}$, remains the same.
Figure 5: Spin density polarization, $m(\bm{r})=\rho_{\uparrow}(\bm{r})-\rho_{\downarrow}(\bm{r})$, of the $10$-$\alpha\alpha$ graphene-fluorographene system. (a) Side view: red and blue lobules represent opposite spin polarizations. (b) Top view: notice the antiferromagnetic order along the transverse direction to the interfaces and ferromagnetic order along the longitudinal direction. The system is globally antiferromagnetic (AF) with higher spin polarization at the channel edges. (c), (d) Integration of the local density of states (ILDOS) within the energy window ($\mathrm{V}$) defined in Fig. 4(c). These states, located below the Fermi energy, exhibit $\mathrm{p_{z}}$ character, with higher weight at the channel edges.
@@ IV.2 $\alpha\beta$ channels
Figure 6: Spin-polarized band structure of $n$-$\alpha\beta$ graphene-fluorographene systems for $n=4,6,8$ and $10$. Red and blue dots represent opposite spin states ($\uparrow$ and $\downarrow$). The ferromagnetic ground state leads to a spin-split band structure around the Fermi level. The band structure also reveals a semiconductor-metal transition with increasing channel width. At $\mathrm{k}=0$, the energy gap between the conduction ($b_{c}$) and valence ($b_{v}$) bands decreases with increasing $n$.
Figure 7: Three-dimensional plot of the spin density polarization, $m(\bm{r})$, of the $n$-$\alpha\beta$ systems for $n=6,8$, and $10$. (a), (e) Top views; (b)–(d) side views. The spin polarization is higher at the channel edges and decreases towards the center, with red and blue lobules representing opposite spin polarizations. The weight at the $\alpha$ edge remains constant, while it weakens at the $\beta$ edge with increasing channel width ($n\geq 8$) and nearly vanishes for (d), (e) $n=10$. Top views (a) and (e) reveal a ferromagnetic order along the longitudinal direction and an antiferromagnetic order along the transverse direction. Despite this, the net magnetization is nonzero, making the $n$-$\alpha\beta$ nanoribbons ferromagnetic.
Figure 8: Integrated local density of states (ILDOS) within the energy window ($\mathrm{V}$) defined in Fig. 6 for $\alpha\beta$ graphene-fluorographene systems. Similar to the $\alpha\alpha$ systems [Figs. 5(c) and 5(d)], the $\alpha\beta$ states near the Fermi level exhibit $\mathrm{p_{z}}$ character, with their weight concentrated along the graphene channel.
@@
A comparison of the band structures of $n$-$\alpha\alpha$ (Fig. 4) and $n$-$\alpha\beta$ (Fig. 6) reveals that the most significant differences, without considering the spin splitting, occur at $\mathrm{k}\sim 0$. This indicates that the states of the valence and conduction bands near $\mathrm{k}\sim 0$ are strongly influenced by the distinct chemical nature of the $\beta$ interface, the key structural difference between the two systems.
@@
Figure 8 shows the ILDOS within the energy window $\mathrm{V}$ marked in Fig. 6. For the $6$-$\alpha\beta$ nanoribbon [Fig. 8(a)], the ILDOS strongly resembles the spatial profile of the spin density polarization [Fig. 7(b)], indicating that the states within the window $\mathrm{V}$ are the primary contributors to the spin-polarized regions. For the wider $8$-$\alpha\beta$ and $10$-$\alpha\beta$ systems, however, a notable difference arises at the $\beta$ edge between the ILDOS [Figs. 8(b) and 8(c)] and the spin density polarization [Figs. 7(c) and 7(d)] due to the reduced spin polarization of that edge, in agreement with the evolution of the band structure near $\mathrm{k}\sim 0$ (Fig. 6) with increasing channel width. Table 2 summarizes the key properties of $n$-$\alpha\beta$ graphene-fluorographene nanoribbons.
@@ V Tight-binding model
Figure 9: Electronic band structure and projected density of states (PDOS) for bulk graphene and fluorographene. (a) Graphene band structure, with the $\pi$ and $\sigma$ states labeled. (b) Graphene PDOS. (c) Fluorographene band structure. (d), (e) Fluorographene PDOS projected onto F and C atoms. States labeled $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{XY}$ in (c) will be used to describe the electronic structure of $n$-$\alpha\alpha$ and $n$-$\alpha\beta$ nanoribbons in the following sections.
@@
To identify the orbital character of the hybridized states in the nanoribbons, we first analyze the electronic structure of bulk graphene and fluorographene, along with their projected densities of states (PDOS), as shown in Fig. 9. Bulk graphene exhibits characteristic $\pi$ states near $\mathrm{E_{F}}$ with $\mathrm{p_{z}}$ orbital character [28]. Additionally, $\sigma$ states lie further away from $\mathrm{E_{F}}$ [Figs. 9(a) and 9(b)]. In fluorographene [Figs. 9(c)–9(e)], we identify three key sets of states: (i) the $\mathcal{A}$ states, which are above $\mathrm{E_{F}}$ and have a dominant $\mathrm{p_{z}}$ character mixed with $\mathrm{s}$ states; (ii) the $\mathcal{XY}$ states, which are below $\mathrm{E_{F}}$ and have primarily $\mathrm{p_{xy}}=\mathrm{p_{x}}+\mathrm{p_{y}}$ character; and (iii) the $\mathcal{B}$ states, which are found below $-4$ eV, and have mixed $\mathrm{p_{z}}$ and $\mathrm{s}$ character.
Figure 10: Band structure of the (a)-(c), (g), (h) $6$-$\alpha\alpha$ and (d)-(f), (i), (j) $6$-$\alpha\beta$ systems. The projection (weight) on some relevant orbitals is highlighted. The size of each ($\mathrm{k},\mathrm{E}$) point reflects the weight of the eigenstates. The bands are classified based on the dominant orbital character from the reference states in Fig. 9: $\pi$ states ($\mathrm{p}_{z}$ character) from graphene (a), (d). $\mathcal{A}$ and $\mathcal{B}$ states ($\mathrm{p}_{z}$ character) mainly located in the fluorographene zones (b), (c), (e), (f). $\mathcal{XY}$ states ($\mathrm{p_{x}+p_{y}}$ character) mainly located in the fluorographene zones (g)–(j). In the $6$-$\alpha\alpha$ system, the conduction band ($b_{c}$) near $\mathrm{k}\sim$0 exhibits a hybridization of $\pi$ and $\mathcal{A}$ states, while the valence band ($b_{v}$) is primarily composed of $\pi$ states. A similar behavior is observed for the $6$-$\alpha\beta$ system, where both $b_{c}$ and $b_{v}$ bands are dominated by $\pi$ states.
@@
    Dominant $\mathrm{\pi}$ character: Figs. 10(a) and 10(d) highlight the dominance of $\mathrm{\pi}$ states ($\mathrm{p_{z}}$ orbitals from graphene carbon atoms) near $\mathrm{E_{F}}$, spanning the energy range [8[-8[ - 8 eV, $12$ eV]]]]. The valence ($b_{v}$) and conduction ($b_{c}$) bands are primarily composed of these $\mathrm{\pi}$ states, but exhibit crossings with bands of other orbital character.
@@
    Hybridization with $\mathcal{A}$ states: Figs. 10(a)–10(f) reveal hybridization between $\mathrm{\pi}$ and $\mathcal{A}$ states contributing to the conduction band, with dominant C-$\mathrm{p_{z}}$ character. This is evident from the states within [1[1[ 1 eV, $12$ eV]]]] exhibiting both $\mathrm{\pi}$ and $\mathcal{A}$ orbital character due to their presence in both the graphene and fluorographene regions. The spin splitting observed in the $n$-$\alpha\beta$ bands near $\mathrm{k}=0$ (particularly in the 22226666 eV window) further supports this $\mathrm{\pi}$-$\mathcal{A}$ hybridization.
@@
    Minimal role of $\mathcal{XY}$ and $\mathcal{B}$ states: The lower panels of Fig. 10 demonstrate negligible hybridization between the $\mathrm{\pi}$ and $\mathcal{XY}$ states (dominant F-$p_{xy}$ character, located within [1[-1[ - 1 eV, $-8$ eV]]]]). The absence of simultaneous $\mathcal{XY}$ and $\mathrm{\pi}$ character within the $b_{v}$ band, along with the lack of spin polarization affecting $\mathcal{XY}$ states, establishes their minimal contribution to $b_{v}$. Similarly, the $\mathcal{B}$ states (dominant F-$\mathrm{p_{z}}$ character, within [4[-4[ - 4 eV, $-11$ eV]]]]) lie well below the valence and conduction bands, making their contribution negligible.
As a consequence, to construct a simplified electronic model reproducing the $b_{v}$ and $b_{c}$ bands, the hybridization between the $\mathrm{\pi}$ and $\mathcal{A}$ states is the crucial ingredient.